OpenStudy (anonymous):

Find the number halfway between 3/5 and 7/10

OpenStudy (imstuck):

What you are doing when you do this is finding the average. To make it simple, add them together and divide the result in half. That is the mid-point between them.

OpenStudy (anonymous):

5/5=1 I guess

OpenStudy (anonymous):

2/3

OpenStudy (imstuck):

13/10 divided by 2?

OpenStudy (anonymous):

I did 10/15 divided by 3

OpenStudy (anonymous):

After you've added the fractions together. \[\frac{ 3 }{ 5 }+\frac{ 7 }{ 10 }=\frac{ 13 }{ 10 }\] You divide by two. Since you are looking for the average. There are two numbers/fractions. \[\frac{ 13 }{ 10 } \div2=answer\]

OpenStudy (imstuck):

Try it as Dscdago did it. He is correct.